# SpikeBottleNet: Energy Efficient Spike Neural Network Partitioning for Feature Compression in Device-Edge Co-Inference Systems

### v1

@@ 4.1 Experimental Setup
For our analysis, we opted for the widely recognized ResNet50 architecture. We consider each residual block (RB) in ResNet as a potential splitting point. The ResNet50 model contains 16 residual blocks, leading to the creation of 16 distinct models. Each model is defined by placing the bottleneck unit after one of the 16 RBs. This placement results in variable degrees of reduction in both the channel and spatial dimensions across 16 different configurations. The structural design of the ResNet50 is illustrated in Figure 2.
@@
Spike ResNet50 4 76.81
2 74.69
Table 2: Accuracy of the Spike ResNet50 model at different time steps
While our spike version of ResNet50 may not surpass the current state-of-the-art performance on the CIFAR-100 dataset, the primary focus of this study lies in assessing the compression capabilities and energy efficiency achieved by our approach. As reported in the Table 2, during the training phase, the experiments achieved an accuracy of 76.81% for the Spike ResNet50 model with a time step of 4 and 74.69% with a time step of 2. To achieve a high feature compression ratio, it is essential to limit the time step to no more than 4. The cross-entropy function served as the loss function during the training process.
@@ 4.2 Compression Capability Comparison
Traditional DNNs based device-edge co-inference systems use continuous values for activations and feature compression during both training and inference phases [6]. These systems typically require high precision, often employing 32-bit floating-point (FP) representations, to preserve model accuracy. However, techniques like quantization [17] can reduce the bit-width requirements, often down to 8 bits or even lower. Conversely, our SNN based SpikeBottleNet relies on spikes, which inherently lower the bit requirements for encoding neuron states. These spikes can be efficiently encoded using just 1-bit representations, which significantly reduces the communication overhead and bitstream size. In the following, we compare the compression capability of our SpikeBottleNet with other methods, such as BottleNet++ across different splitting points.
Figure 3: Compression Capability Comparison between SpikeBottleNet and BottleNet++
@@
The results depicted in Figure 3 reveal that, across all splitting points SpikeBottleNet consistently achieves the lowest communication overhead. This efficiency is highlighted by its achievement of up to 256x compression ratio in the final convolutional layer of the Spike ResNet50 architecture, compressing 131,072 bits (16 KB) down to 512 bits (64 Bytes), with an optimal accuracy reduction of only 2.52%. In comparison, SpikeBottleNet demonstrates a 50% reduction in the data size of 64 bytes required by the BottleNet++ (128 bytes).
A key advantage of our proposed model is its ability to facilitate earlier layer partitioning within the SNN architecture. In earlier layers, such as RB2 and RB3, SpikeBottleNet achieves a remarkable bit reduction of up to 75% compared to BottleNet++. This enhanced feature compression capability extends to layers RB9, RB10, RB11, RB12 and RB14 as well. Importantly, with the exception of RB1, SpikeBottleNet achieves at least a 50% reduction in data size across all evaluated splitting points compare to the baseline. This comparison emphasizes the SpikeBottleNet’s superior capability in feature compression for efficient communication than traditional compressors.
We study the trade-off between feature map compression and its impact on model accuracy. The Table 1 illustrates the model accuracy drop, the size of compressed features, and the bit reduction at various splitting points, alongside the time steps employed for these points. Our analysis reveals that accuracy loss remains below 3% for most of the splitting points. In some cases, however, we relaxed the accuracy threshold to 4% (RB1, RB6, RB9, RB10, and RB14) to achieve even more significant reduction in bit ratio. Notably, at RB8, no configuration could be identified that compressed features while maintaining accuracy loss of below 3%. Another key observation we noticed in our experiment is that the spatial reduction only worked for the final convolutional layer at RB16. At all other partitioning points, only channel reduction maintained the desired accuracy degradation threshold.
A key objective was to preserve a time step of 2 for compression across all layers to maximize the compression ratio. However, our findings suggest that time step 2 is only effective mostly for earlier layers, up to RB11, without a significant impact on accuracy. Beyond this point, increasing the time step to 4 becomes necessary. This suggests that the density of useful information in the feature maps increase in deeper layers. While initial layers have more redundancy that can be effectively compressed, deeper layers encode more critical information in more compact forms, thus requiring finer temporal resolution (using time step 4) to preserve this information during compression.
32 bit-FP [10]: MAC 4.6pJ, AC 0.9 pJ
Splitting Point RB1 RB2 RB3 RB4 RB5 RB6 RB7 RB8 RB9 RB10 RB11 RB12 RB13 RB14 RB15 RB16
BottleNet++ $(\mathrm{m}J)$ 0.27 0.60 0.93 1.33 1.66 1.99 2.32 2.72 3.05 3.38 3.70 4.03 4.36 4.77 5.09 5.42
SpikeBottleNet $(\mathrm{mJ})$ 0.15 0.10 0.15 0.23 0.26 0.29 0.31 0.72 0.39 0.42 0.46 0.50 0.99 0.57 1.13 1.05
$\frac{E(\text{ BottleNet }++)}{E(\text{ SpikeBottleNet })}$ 1.84 6.09 6.17 5.82 6.34 6.84 7.40 3.76 7.77 7.95 8.00 8.11 4.42 8.42 4.52 5.18
ROLLS chip[14]: 77fJ per SyOPs
SpikeBottleNet $(\mathrm{m}J)$ 0.013 0.008 0.013 0.020 0.022 0.025 0.027 0.062 0.034 0.036 0.040 0.043 0.084 0.048 0.096 0.090
$\frac{E(\text{ BottleNet }++)}{E(\text{ SpikeBottleNet })}$ 21.50 71.17 72.06 68.04 74.06 79.95 86.50 43.93 90.85 92.88 93.49 94.80 51.64 98.42 52.83 60.54
Table 3: Comparison of Mobile Energy Consumption Across Various Partition Points for Traditional 45nm Hardware and ROLLS Neuromorphic Hardware
@@ 4.3 Theoretical Energy Efficiency Estimation
Our analysis of SpikeBottleNet’s theoretical energy consumption is based on the established research [12][11]. We assume an implementation on 45nm mobile hardware, where the energy cost per MAC operation is 4.6pJ and per AC operation is 0.9pJ [10]. Our calculated energy consumption represents the edge device’s energy usage up to a specific residual block (splitting point). This analysis concentrates on the convolutional (CONV) layers within the residual blocks, which constitute a significant portion of the network’s FLOPs. The initial CONV layer converting images to spikes and the down sampling shortcuts are excluded as they deviate from the criteria for spike-based convolutions. Similarly, computations from the edge device’s encoder are omitted due to their constant operational cost.
@@
The firing rate $fr$ is defined as the probability of a neuron firing per timestep. It is estimated under a batch of random samples using the formula:
$fr\ =\ \frac{\#Spikes}{\#Neurons}\ \ast\ T$ (8)
where #$\#Spike$ and #$\#Neuron$ respectively denote the total spikes in $T$ time steps and the number of neurons in the network.
The results in Table 3 demonstrate that, on 45nm hardware, SpikeBottleNet exhibits greater power efficiency compared to BottleNet++ across all splitting points. Specifically, it achieves over 8 times at some splitting points (RB11, RB12, and RB14).
@@
To estimate energy consumption on neuromorphic chip, we considered the ROLLS neuromorphic chip. Fabricated using a 180nm CMOS process, the ROLLS chip reportedly consumes 77fJ per SyOPs [14]. As reported in the Table 3, SpikeBottleNet exhibits up to 98 times greater power efficiency at RB14 on ROLLS hardware than BottleNet++ on conventional hardware. This finding underscores the promising future of SNN-based SpikeBottleNet. The estimated power consumption of SNNs implemented on neuromorphic hardware significantly surpasses that of DNN-based BottleNet++.
Pretraining a SNN, including the SpikeBottleNet architecutre, remains an energy efficiency challenge. The extended training time, often lasting a day even on powerful GPUs, poses a significant challenge. To address this challenge, the further development of novel algorithms and specialized hardware accelerators specifically designed for SNNs is necessary. These innovations hold the key to significantly reducing training costs and paving the way for a new era of energy-efficient neuromorphic computing.
5 Conclusions and Future Work
@@
This work presents a novel collaborative intelligence framework, SpikeBottleNet, that tackles the latency and energy limitations hindering the deployment of deep neural networks on mobile devices. Our Spiking Neural Network (SNN) partitioning strategy significantly reduces the communication cost of transferring features from the mobile device to the cloud, while concurrently minimizing the mobile device’s energy consumption during co-inference operations at the device-edge. By exploiting the inherent binary spike signals employed by SNNs, SpikeBottleNet utilizes low-power accumulation methods instead of the high-power multiply-accumulate operations. This substitution leads to significantly minimizing the mobile device’s energy footprint. Additionally, SpikeBottleNet allows for earlier network partitioning, achieving higher compression rates than those attained by current methods. Our evaluations demonstrate that the proposed SNN-based SpikeBottleNet achieves a remarkable 256x compression ratio in the final convolutional layer. Furthermore, it translates to a concurrent 75% reduction in bit transmission at some of the earlier splitting points compared to the baseline model, with a measured accuracy degradation of only less than 3%.
Future research will focus on two key areas: communication channel modeling and input feature encoding. We aim to explore how channel characteristics like noise and bandwidth limitations impact SpikeBottleNet’s performance. This will involve developing adaptive encoding and error-correction techniques for robust spike signal transmission under varying network conditions. Additionally, we plan to investigate various input encoding methods, including attention-based mechanisms, to further enhance the network’s efficiency and adaptability.

### v2

@@ 4.1 Experimental Setup
For our analysis, we opted for the widely recognized ResNet50 and MobileNetV1 architectures. We identified each residual block (RB) within ResNet50 and each inverted residual block (IRB) within MobileNetV1 as potential points for architectural modification. The ResNet50 model comprises 16 RBs, while the MobileNetV1 architecture contains 13 IRBs, resulting in the creation of an equal number of distinct models. Each model is defined by strategically positioning the bottleneck unit after one of these blocks. This placement yields varying degrees of reduction in both the channel and spatial dimensions across the respective configurations.
Architecture Split Pt. T-Step Acc. (Drop) Bit Red. Energy Imp.
ResNet50 4 2 73.31 (0.76)16x 10x
4 75.16 (1.1)32x 6x
MobileNetv1 3 2 69.82 (1.21)16x 8x
4 73.37 (1.47) % 64x 4x
Table 1: Impact of Time Steps on SpikeBottleNet Performance
@@
ResNet50 4 76.27
2 74.07
MobileNetV1 4 74.86
2 71.03
Table 2: Accuracy of the spike version of the models at different time steps on CIFAR100 dataset
While our spike version of ResNet50 and MobileNetV1 may not surpass the current state-of-the-art performance on the CIFAR-100 dataset, the primary focus of this study lies in assessing the performance degradation, compression capabilities and energy efficiency achieved by our approach. The training phase accuracies of the models are reported in the Table 2. To maintain a high feature compression ratio, the time step should be limited to 4 or fewer, as additional time steps linearly increase latency and energy consumption. The cross-entropy function served as the loss function during the training process.
@@ 4.2 Compression Capability Comparison
Traditional DNNs based edge-cloud co-inference systems use continuous values for activations and feature compression during both training and inference phases [3]. These systems typically require high precision, often employing 32-bit floating-point (FP) representations, to preserve model accuracy. However, techniques like quantization [30] can reduce the bit-width requirements, typically from 32-bit floating-point (FP32) to lower bit-width formats like 16-bit (FP16), 8-bit (INT8), or even 4-bit. However, the most significant challenge with reducing bit width is the potential loss in model accuracy. Furthermore, lower precision formats can lead to rounding errors, reduced precision in gradient updates, and less representational capacity. Conversely, our SNN based SpikeBottleNet relies on spikes, which inherently lower the bit requirements for encoding neuron states. These spikes can be efficiently encoded using just 1-bit representations, which significantly reduces the communication overhead and bitstream size. In the following, we compare the compression capability of our SpikeBottleNet with other methods, such as BottleNet [3] [4] across different splitting points. The BottleNet architecture used in this comparison is our implemented version.
Figure 2: Compression capability comparison between SpikeBottleNet and BottleNet
@@
The results depicted in Figure 2 reveal that, across all splitting points with a time step 2, SpikeBottleNet consistently achieves the lowest communication overhead compared to the DNN-based BottleNet. This efficiency is highlighted by its achievement of up to 256x compression ratio in the final convolutional layer of the SpikeBottleNet + ResNet50 architecture, compressing 65,536 bits (8 KB) down to 256 bits (32 Bytes), with an optimal accuracy reduction of only 0.16%. In comparison, SpikeBottleNet demonstrates a 75% reduction in the data size of 32 bytes required by the BottleNet + ResNet50 (128 bytes). This 75% reduction in data size by SpikeBottleNet occurs consistently across all evaluated splitting points compared to the baseline. Similarly, the compression strategies for SpikeBottleNet + MobileNet effectively reduce data dimensions, achieving significant compression ratios, particularly at final layers (up to 174x), with minimal accuracy loss of 0.720%. This highlights SpikeBottleNet’s superior feature compression capability, enabling more efficient communication than traditional compression methods.
@@ 4.3 Theoretical Energy Efficiency Estimation
Our analysis of SpikeBottleNet’s theoretical energy consumption is based on the established research [15][31]. We assume an implementation on 45nm mobile hardware, where the energy cost per MAC operation is 4.6pJ and per AC operation is 0.9pJ [32]. Since SpikeBottleNet operates without requiring MAC operations and relies solely on AC, its significantly lower energy consumption is evident. Our calculated energy consumption represents the edge device’s energy usage up to a specific splitting point. This analysis concentrates on the convolutional (CONV) layers within the residual blocks, which constitute a significant portion of the network’s FLOPs. The computational cost from the edge device’s encoder is excluded from analysis, as the encoder is a lightweight layer and its operational cost remains constant.
@@
The firing rate $fr$ is defined as the probability of a neuron firing per timestep. It is estimated under a batch of random samples as follows:
$fr=\frac{1}{LN}\sum_{l=1}^{L}\sum_{i=1}^{N}o_{i,l}$ (8)
where $L$ is the total number of layers until splitting points, $N$ is the number of neurons per layer, and $o_{i,l}$ represents the spike activity of neuron $i$ in layer $l$.
Figure 3: Comparison of mobile energy consumption across various partition points for traditional 45nm hardware and ROLLS neuromorphic hardware
The results in Figure 3 demonstrate that, on both traditional and neuromorphic hardware, SpikeBottleNet exhibits greater power efficiency compared to BottleNet across all splitting points. For instance, it achieves over 12x energy reduction on 45nm hardware at the first splitting point of the ResNet architecture. Similary, SpikeBottleNet offers a 10x less energy consumption compared to BottleNet at the first splitting point of MobileNetV1. This trend holds consistently across most of the splitting points.
@@
To estimate energy consumption on neuromorphic chip, we considered the ROLLS neuromorphic chip. Fabricated using a 180nm CMOS process, the ROLLS chip reportedly consumes 77fJ per SyOPs [20]. In our experiments, SpikeBottleNet exhibits up to 144x greater power efficiency at first splitting points of ResNet on ROLLS hardware compare to BottleNet on conventional hardware. Detailed results are provided in Supplementary Materials Section 8. This finding underscores the promising future of SNN-based SpikeBottleNet. The estimated power consumption of SNNs implemented on neuromorphic hardware significantly surpasses that of DNN-based BottleNet.
4.4 Ablation Study
4.4.1 Impact of Time Step Changes on SpikeBottleNet’s Performance
This ablation study investigates the impact of varying time steps within the SpikeBottleNet architecture on accuracy, bit reduction, and energy efficiency when deployed on conventional 45nm hardware. We selected one-fourth of the splitting points in each network for the ablation study. By dividing the network into four equal segments, we minimize computational demands on edge devices and reduce power consumption. Our experiments reported in Table 1 demonstrate that increasing time steps effectively enhances bit reduction but introduces a trade-off in energy efficiency. For example, with 4 time steps, ResNet50 achieves a 32x reduction in bitrate and a 6x improvement in energy consumption, while MobileNetv1 realizes a 64x bit reduction with a 4x energy gain. Detailed results are available in the supplementary materials Section 6 and Section 7.
While increasing time steps generally leads to improved accuracy and data compression, it is essential to carefully consider the potential trade-offs in energy efficiency and latency when deploying SpikeBottleNet in edge-cloud environments.
5 Conclusions
@@
This work presents a novel collaborative intelligence framework, SpikeBottleNet, that tackles the energy limitations hindering the deployment of deep neural networks on mobile devices. Our SNN partitioning strategy significantly reduces the communication cost of transferring features from the mobile device to the cloud, while minimizing the mobile device’s energy consumption during co-inference at the edge-cloud. By exploiting the inherent binary spike signals employed by SNNs, SpikeBottleNet utilizes low-power accumulation methods instead of the high-power multiply-accumulate operations. This substitution leads to significantly minimizing the mobile device’s energy footprint. Our evaluations demonstrate that the proposed SNN-based SpikeBottleNet achieves a remarkable 256x bit compression ratio in the final convolutional layer. Moreover, it translates to a 75% reduction in bit transmission at most of the splitting points compared to the baseline model, with a measured accuracy degradation of only less than 2%. Furthermore, SpikeBottleNet is well-suited for deployment on neuromorphic hardware, offering up to 144x energy savings on mobile devices.
Future work will focus on reducing SpikeBottleNet’s timesteps from multiple to a single step, aiming to further minimize latency, communication costs, and energy usage. Additional goals include exploring dynamic timesteps to potentially enhance accuracy, compression, and energy efficiency, optimizing the model for even more effective edge-cloud co-inference solution.
Supplementary Material
6 Impact Analysis of Compression Techniques on ResNet50
The Table 3 and Table 4 provide a detailed analysis of the impact of applying BottleNet and SpikeBottleNet compression techniques to ResNet50 at different network layers. The analysis considers feature dimensions, compressed feature dimensions, required byte count, compression ratios, and accuracy loss. As compression is applied to deeper layers, feature dimensions diminish, and compression ratios generally increase, reaching a maximum of 256x at layer 16. While higher compression ratios often result in substantial accuracy drops, as exemplified by the 1.89% decrease at layer 8, some deeper layers, such as layer 16, achieve high compression (256x) with minimal accuracy loss (0.16%). This analysis underscores the delicate balance between data reduction and model performance across various network depths.
7 Impact Analysis of Compression Techniques on MobileNetV1
Table 5 and 6 present the performance of splitting SpikeBottleNet at different layers of MobileNetV1. Early layer splits (1-3) result in significant data size reductions (25,088 bytes) with moderate compression ratios (16-32x). Notably, these early splits incur only minor accuracy drops ( 1.3%). As the splitting point moves deeper into the network (layers 12-13), the compression ratios increase dramatically (up to 174x), reducing data sizes to as low as 288 bytes. However, the impact on accuracy becomes more pronounced, with drops ranging from 0.72% to 2.04%. This analysis underscores a clear trade-off between compression and accuracy. Deeper splits offer higher compression but may compromise model performance, whereas earlier splits provide a more balanced approach, trading off some compression for improved accuracy.
Splitting Point 1 2 3 4 5 6 7 8
Time Step 2 2 2 2 2 2 2 2
Original 256x32x32 256x32x32 256x32x32 512x16x16 512x16x16 512x16x16 512x16x16 1024x8x8
Compressed 16x32x32 16x32x32 16x32x32 32x16x16 16x16x16 16x16x16 16x16x16 128x8x8
BottleNet (B) 16,384 16,384 16,384 8,192 4,096 4,096 4,096 8,192
SpikeBottleNet (B) 4,096 4,096 4,096 2,048 1,024 1,024 1,024 2,048
Compression Ratio 16 16 16 16 32 32 32 8
Accuracy Drop (%) 0.67 0.27 0.09 0.36 0.98 0.36 0.21 1.89
Table 3: Compression and Accuracy Trade-offs Across Different Splitting Points on ResNet50 architecture (1 - 8) in SpikeBottleNet
Splitting Point 9 10 11 12 13 14 15 16
Time Step 2 2 2 2 2 2 2 2
Original 1024x8x8 1024x8x8 1024x8x8 1024x8x8 1024x8x8 2048x4x4 2048x4x4 2048x4x4
Compressed 32x8x8 16x8x8 16x8x8 16x8x8 16x8x8 16x4x4 16x4x4 8x4x4
BottleNet (B) 2,048 1,024 1,024 1,024 1,024 256 256 128
SpikeBottleNet (B) 512 256 256 256 256 64 64 32
Compression Ratio 32 64 64 64 64 128 128 256
Accuracy Drop (%) 0.83 1.60 1.36 1.23 1.06 1.53 1.22 0.16
Table 4: Extended Compression and Accuracy Trade-offs Across Different Splitting Points on ResNet50 architecture (8 - 16) in SpikeBottleNet
Splitting Point 1 2 3 4 5 6 7
Time Step 2 2 2 2 2 2 2
Original 64x112x112 128x56x56 128x56x56 256x28x28 256x28x28 512x14x14 512x14x14
Compressed 8x56x56 8x56x56 8x56x56 16x28x28 16x28x28 32x14x14 32x14x14
BottleNet (B) 25,088 25,088 25,088 12,544 12,544 6,272 6,272
SpikeBottleNet (B) 6,272 6,272 6,272 3,136 3,136 1,568 1,568
Compression Ratio 32 16 16 16 16 16 16
Accuracy Drop (%) 1.34 1.35 1.21 1.47 0.78 0.77 0.73
Table 5: Compression and Accuracy Trade-offs Across Different Splitting Points on MobileNetV1 architecture (1 - 7) in SpikeBottleNet
Splitting Point 8 9 10 11 12 13
Time Step 2 2 2 2 2 2
Original 512x14x14 512x14x14 512x14x14 512x14x14 1024x7x7 1024x7x7
Compressed 32x14x14 32x14x14 32x14x14 32x14x14 32x7x7 32x3x3
BottleNet (B) 6,272 6,272 6,272 6,272 1,568 288
SpikeBottleNet (B) 1,568 1,568 1,568 1,568 392 72
Compression Ratio 16 16 16 16 32 174
Accuracy Drop (%) 0.58 0.52 0.84 1.16 2.04 0.72
Table 6: Extended Compression and Accuracy Trade-offs Across Different Splitting Points on MobileNetV1 architecture (8 - 13) in SpikeBottleNet
8 Impact Analysis of Energy Efficiency of SpikeBottleNet and BottleNet Architectures
The provided Tables 7 - 10 compare the energy consumption of SpikeBottleNet and BottleNet architectures when applied to ResNet50 and MobileNetV1 models at various splitting points. For both architectures, SpikeBottleNet consistently outperforms BottleNet in terms of energy efficiency.
For ResNet50, the energy efficiency ratio stabilizes at approximately 10-11 for the 45nm chip and 120-130 for the ROLLS chip at higher splitting points. In contrast, MobileNetV1 displays slightly lower stabilization values, ranging from 7-9 for the 45nm chip and 88-102 for the ROLLS chip. This suggests that SpikeBottleNet’s energy efficiency advantage is more pronounced in MobileNetV1, particularly at lower splitting points.
32 bit-FP: MAC 4.6pJ, AC 0.9 pJ
Splitting Point 1 2 3 4 5 6 7 8
GFLOPs 0.08 0.15 0.22 0.34 0.41 0.49 0.56 0.68
BottleNet (mJ) 0.36 0.69 1.02 1.58 1.91 2.24 2.57 3.13
SpikeBottleNet 45nm (mJ) 0.03 0.06 0.09 0.15 0.19 0.22 0.26 0.33
$\frac{E(\text{ BottleNet })}{E(\text{ SpikeBottleNet })}$ 12.37 11.88 11.56 10.35 10.16 10.05 9.91 9.59
ROLLS chip: 77fJ per SyOPs
Firing Rate 0.2066 0.2151 0.2210 0.2468 0.2515 0.2542 0.2579 0.2664
GSyOPs 0.03 0.06 0.10 0.17 0.21 0.25 0.29 0.36
SpikeBottleNet ROLLS (mJ) 0.002 0.005 0.008 0.013 0.016 0.019 0.022 0.028
$\frac{E(\text{ BottleNet })}{E(\text{ SpikeBottleNet })}$ 144.58 138.87 135.16 121.03 118.77 117.51 115.82 112.13
Table 7: Energy Consumption Comparison for Splitting Points 1 to 8 of ResNet50
32 bit-FP: MAC 4.6pJ, AC 0.9 pJ
Splitting Point 9 10 11 12 13 14 15 16
GFLOPs 0.75 0.82 0.89 0.97 1.04 1.16 1.23 1.30
BottleNet (mJ) 3.46 3.78 4.11 4.44 4.77 5.33 5.66 5.99
SpikeBottleNet 45nm (mJ) 0.35 0.37 0.40 0.42 0.46 0.50 0.52 0.55
$\frac{E(\text{ BottleNet })}{E(\text{ SpikeBottleNet })}$ 9.83 10.17 10.34 10.47 10.47 10.58 10.87 10.87
ROLLS chip: 77fJ per SyOPs
Firing Rate 0.2599 0.2513 0.2471 0.2442 0.2440 0.2416 0.2350 0.2352
GSyOPs 0.39 0.41 0.44 0.47 0.51 0.56 0.58 0.61
SpikeBottleNet ROLLS (mJ) 0.030 0.032 0.034 0.036 0.039 0.043 0.045 0.047
$\frac{E(\text{ BottleNet })}{E(\text{ SpikeBottleNet })}$ 114.93 118.86 120.88 122.32 122.42 123.63 127.11 127.00
Table 8: Extended Energy Consumption Comparison for Splitting Points 9 to 16 of ResNet50
32 bit-FP: MAC 4.6pJ, AC 0.9 pJ
Splitting Point 1 2 3 4 5 6 7
GFLOPs 0.04 0.07 0.13 0.15 0.21 0.23 0.28
BottleNet (mJ) 0.19 0.32 0.58 0.70 0.95 1.07 1.31
SpikeBottleNet 45nm (mJ) 0.02 0.04 0.07 0.09 0.12 0.14 0.17
E(BottleNet) / E(SpikeBottleNet) 11.28 8.35 8.12 7.53 7.66 7.55 7.88
ROLLS chip: 77fJ per SyOPs
GSyOPs 0.02 0.04 0.08 0.10 0.14 0.16 0.18
Firing Rate 0.2265 0.3059 0.3146 0.3396 0.3337 0.3384 0.3242
SpikeBottleNet ROLLS (mJ) 0.001 0.003 0.006 0.008 0.011 0.012 0.014
$\frac{E(\text{ BottleNet })}{E(\text{ SpikeBottleNet })}$ 131.88 97.65 94.95 87.96 89.51 88.27 92.13
Table 9: Energy Consumption Comparison for Splitting Points 1 to 7 of MobileNetV1
32 bit-FP: MAC 4.6pJ, AC 0.9 pJ
Splitting Point 8 9 10 11 12 13
GFLOPs 0.34 0.39 0.44 0.49 0.52 0.57
BottleNet (mJ) 1.55 1.79 2.03 2.28 2.40 2.63
SpikeBottleNet 45nm (mJ) 0.19 0.22 0.24 0.27 0.28 0.30
$\frac{E(\text{ BottleNet })}{E(\text{ SpikeBottleNet })}$ 8.23 8.30 8.37 8.41 8.53 8.76
ROLLS chip: 77fJ per SyOPs
GSyOPs 0.21 0.24 0.27 0.30 0.31 0.33
Firing Rate 0.3105 0.3078 0.3054 0.3038 0.2997 0.2918
SpikeBottleNet ROLLS (mJ) 0.016 0.018 0.021 0.023 0.024 0.026
$\frac{E(\text{ BottleNet })}{E(\text{ SpikeBottleNet })}$ 96.20 97.04 97.81 98.32 99.67 102.37
Table 10: Energy Consumption Comparison for Splitting Points 8 to 13 of MobileNetV1